OpenStudy (aravindg):

Find domain and range of function f(x)=sqrt2-x+sqrt(1+x)

OpenStudy (aravindg):

...........

OpenStudy (aravindg):

domain and range plzz

OpenStudy (aravindg):

reply plzzzzzz

OpenStudy (anonymous):

Roots must be zero or positive. So at first 2 is > than 0 so ok. In the second root it must be \[1+x \ge0 =x \ge-1\] So the domain of the function is \[f(x)\in[-1,+\infty)\] As for the range of the function it is (sqrt2 +1 till +infinity)

OpenStudy (aravindg):

plz answe my othr question also

OpenStudy (anonymous):

\[{-1, \infty}\]